OpenStudy (anonymous):

Eliminate the parameter. x = 5 cos t, y = 5 sin t

OpenStudy (anonymous):

@Chooch146 @some_someone

OpenStudy (anonymous):

@FLVS-Math

OpenStudy (jdoe0001):

$$ \text{one can say that}\\ x=5cos(t) \implies cos(t)=\cfrac{x}{5}\\ \text{and that}\\ y=5sin(t) \implies sin(t)=\cfrac{y}{5}\\ \text{now, using the pythagorean identity}\\ \color{blue}{sin^2(t)+cos^2(t)=1}\\ $$ one can say that ? can you take it from there?

OpenStudy (anonymous):

how do i get rid of the (t) tho? @jdoe0001

OpenStudy (anonymous):

@Mertsj @timo86m

OpenStudy (anonymous):

@Chooch146

OpenStudy (anonymous):

\[x=5\cos t\\ y=5\sin t\] Like @jdoe0001 said, you have to make use of the Pythagorean identity, \(\sin^2\theta+\cos^2\theta=1\). If you plug in the given expressions for \(x\) and \(y\), you have something like \[(5\sin t)^2+(5\cos t)^2=1\\ 25\sin^2t+25\cos^2t=1\\ 25\left(\sin^2t+\cos^2t\right)=1\\ 25=1\] However, this isn't true, so you have to adjust a bit. Multiply the right side by 25 to make it true: \[(5\sin t)^2+(5\cos t)^2=25\] Finally, to eliminate the parameter \(t\), substitute the parametric equations: \[x^2+y^2=25\]